Determine if the formula T=2π√(m/f/x ) is dimensionally correct

1
Expert's answer
2022-05-13T03:24:28-0400

"F=kx,\implies k=\frac Fx,"

"T=2\pi \sqrt{\frac m{\frac Fx}}=2\pi \sqrt\frac mk," correct.